OpenStudy (anonymous):

how many ways can you get a 9 when rolling 2 dice?

OpenStudy (campbell_st):

it depends on the dice...number of faces... and what is on the faces

OpenStudy (anonymous):

both of the dice are 6 faced dice

OpenStudy (campbell_st):

ok... so draw a table  and just complete it... you'll find all the 9s

OpenStudy (anonymous):

i did what you told me and i got 5 outcomes.

OpenStudy (campbell_st):

 i think you have 1 to many

OpenStudy (anonymous):

oh i see i caught my mistake i added 8 and 1

OpenStudy (campbell_st):

no problem... diagrams and pictures are great tools when doing maths..... they help to show whats happening.

OpenStudy (anonymous):

one more question .. so they asked me to find the possible out comes to get a 6 when rolling a 6 faced dice. well i got 5 outcomes . then they ask what the complement is ? what does that mean

OpenStudy (campbell_st):

the complement means not rolling a 6 P(not 6) = 1 - P(6) so P(6) = 1/6 now applying that to the complement P(not 6) = 1 - 1/6 so the answer is 5/6

OpenStudy (anonymous):

oh okay thanks :)